A curve has an equation of y=sin x. If the area of the curve



y=sin x from x=0 to X=pi is revolved about the y-axis, what is the



volume generated?

1
Expert's answer
2022-06-07T07:45:19-0400
"V=\displaystyle\int_{0}^{\pi}2\pi x\sin xdx"

"\int x\sin xdx=-x\cos x+\int \cos x dx"

"=-x\cos x+\sin x +C"

"V=\displaystyle\int_{0}^{\pi}2\pi x\sin xdx"

"=[2\pi(-x\cos x+\sin x)]\begin{matrix}
 \pi\\
 0
\end{matrix}"

"=2\pi (\pi+0-0)=2\pi^2({units}^3)"